Hi I have to write two equations using two integers between -12 and 12. I have chosen -3 and 2 as my numbers. How do I write the equations? Please advise.
Answer by Theo(13342)

y = -3x + 2
y = 2x - 3
you can solve these equations to find the value of x that will satisfy both of them at the same time.
the value of x that will do that is x = 1
if you replace x with 1, then the equations become:
y = -3*1 + 2 = -3 + 2 = -1
y = 2*1 - 3 = 2 - 3 = -1
with the value of x = 1, y equals -1 in both equations.
this is just one of many equations that can be written with those 2 numbers.